The average age of three boys is 25 years and their ages are in the proportion 3:5:7.The age of the youngest boy is
a)21 b)18 c)15 d)9

Answer by checkley77(12844)

3:5:7
3*25=75 total age of the 3 boys.
3+5+7=15
75/15=5 multiplier for each of the factors 3,5 & 7
3*5=3*5=15 age of the youngest boy.